# Denoising Neural Reranker for Recommender Systems

### v2

@@ 4.4 Sensitivity Analysis (RQ6)
We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3. While the Beta noise generally performs better than Gaussian noise in most settings, as shown in Table 2, there exist some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the synthetic noisy scores in Eq.(4), $\lambda_{m}$ for the combined denoising loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic noise generator to model-based one; 2) Over-amplifying these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components. Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise generator can further enhance the performance over the heuristic one.
@@ 5 Conclusion
All the results in this work are reproducible. We have discussed the optimal hyperparameters and the details on devices and software environments in Section 4.1 and Appendix D.2.
@@ B.1 Derivation of the Three Objectives in our Solution
In this section, we illustrate the derivation of the Eq.(5). Similar to the variational Bayesian inference Kingma & Welling (2014), we start by analyzing the KL divergence between the real retriever scores and synthetic noisy scores:
@@
Note that a key difference between Eq.(8) and the conventional derivation in Kingma & Welling (2014) is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise generator that mimics the behavior of retriever scores. In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the real retriever scores $\mathbf{x}_{u}$ are not sampled by a manually defined distribution. Instead, an existing retriever $p_{\text{x}}$ serves as the prior distribution.
@@ D.3 Baselines
  • SASRec Kang & McAuley (2018) proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
  • Caser Tang & Wang (2018) introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
@@ E.2 Visualization of Learned Noises
Figure 5: The visualization of generaed noise from different variants.
@@ E.5 Experiments on Larger Reranker Backbones
DNR is a backbone-agnostic framework for reranking tasks, utilizing and denoising the noisy scores from the previous retrieval stage through adversarial learning. Recently, LLMs have become a powerful tool for reranking tasks. However, text-based LLM4Rerank is not comparable in our setting offline or online, due to the lack of rich textual information (such as detailed item descriptions) that LLM-based rerankers fundamentally rely on. To evaluate the effectiveness of DNR in large models, we scaled up our PRM baseline by significantly increasing the number of its Transformer layers, serving as a “larger model” for comparison. We then tested our DNR framework, using this scaled-up PRM as its reranker backbone. The results are presented in the table 10.
@@ E.6 Online A/B Test
Detailed Experimental Settings for Online Test The multi-stage pipeline of online system consists of three stages: retrieval, ranking, and reranking. In the final reranking stage, the top-6 items for user exposure are selected from a candidate set of 50 items generated by the ranking stage. During the ranking process, a scoring model computes multiple dimensional features for each item. Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$, which corresponds to the “realshow” metric in the online system. Note that the final reranking stage considers the combination of the retrieval stage and the middle ranking stage as a holistic previous stage, so the output PCTR scores from the ranking stage are considered as “retrieval score” in this final stage. DNR is applied to generate a new set of scores for 50 selected items and the top-6 items are chosen as the final exposure.
Experiment Results For our online experiments, the total traffic is randomly divided, with 12.5% allocated to the DNR method and 12.5% to the baseline method. Each experiment is conducted for at least 7 days to ensure the reliability of the results. The DNR method demonstrates a significant improvement in realshow(realshow refers to the cumulative number of videos watched by users.), increasing it by 1.089% . These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow. Additionally, the industrial recommendation environment involves various other metrics in addition to the realshow metric, including but not limited to app usage time, watch time, share-rate, like-rate, and comment-rate. We provide the corresponding results in Table 11. We can see that, in our scenario, watch-time and share-rate are slightly negatively impacted, while like-rate and comment-rate are slightly positively impacted, indicating a potential trade-off between metrics. In general, the impact on all other metrics is not statistically significant, indicating the effectiveness of our method.
@@
E.7 Computational Analysis
Theoretically, our method does not change the computational cost during inference since we can adopt the same reranking model structure (e.g., PRM Pei et al. (2019)). In the training phase, we include an additional noise generator that may require O(1) computational overhead, whose cost depends on the generator’s structure. Empirically, the generator does not have to be sophisticated, and the computational overhead can be trivial compared to the reranker’s training.
To better demonstrate computational cost, we compare the per-epoch training time of DNR against several baseline rerankers, including PRM Pei et al. (2019), Pier Shi et al. (2023), DCDR Lin et al. (2024), and a "w/ score" variant—where retriever scores are directly used as input to the reranker. In table 12, we present the time cost comparison for different reranker baselines in each training epoch. The variant “w/ score” denotes methods incorporating retriever scores as input features, regardless of the specific integration method (e.g., concatenation, plus, or weighted combination). Our proposed DNR demonstrates competitive training efficiency, with time costs comparable to both PRM and w/ score baselines across all datasets (ML-1M, KuaiVideo, and Book). This validates the computational efficiency of our approach while maintaining competitive performance.
Table 12: Comparison of time costs.
Datasets ML-1M Kuaivideo Book
PRM 01m01s 02m21s 01m42s
Pier 01m54s 04m11s 02m36s
DCDR 02m27s 04m58s 03m02s
w/ score 01m03s 02m49s 01m46s
DNR 01m06s 03m12s 01m51s
Appendix F Limitation and Discussion
F.1 Limitations
Computational Cost: Compared to a direct optimization framework with Eq.(2) that only learns $\theta$, our framework requires extra efforts to train a posterior model $\phi$ (if using learnable posterior) with three additional loss terms to back-propagate. This might multiply the training cost for several folds. Fortunately, $\phi$ is not involved in inference, so the inference cost stays the same.
F.2 Future Directions
Multi-task Scenarios: When a user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$. Yet, different behavior signals may typically follow a heterogeneous distribution where a uniform modeling technique might become sub-optimal. Additionally, the problem may become even more sophisticated when there exist multiple retrievers that estimate the same metric, especially in industrial settings. In this case, one may have to come up with a more articulated integration method to clarify the connections between the user feedback and different retriever scores. Thus, we believe it is worth further investigating how the proposed framework may accommodate the multi-task solutions.
Multi-stage vs. End-to-End: Recent advances in end-to-end recommenders have witnessed significant advances, and evidence Deng et al. (2025) has shown that end-to-end recommenders are potentially more effective since they no longer need to worry about the consistency challenge in the multi-stage framework. Yet, there is still no evidence of a general superiority comparing these two paradigms, and there is still no optimal solution for a fully consistent multi-stage recommender.
Appendix G Use of LLMs
Large language models (LLMs) were used only to aid writing polish—including refining sentence phrasing, logical flow, and prose clarity—without altering original meanings or technical details. LLMs did not participate in core research tasks (e.g., experiment design, data processing, model training, result analysis, or drafting key technical content).

### v3

@@ 4.4 Sensitivity Analysis (RQ6)
We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3. When adjusting $\lambda_{m}$, $\lambda_{e}$ is fixed at 40; when adjusting $\lambda_{e}$, $\lambda_{m}$is fixed at 0.3. For $\lambda_{c}$ adjustments, $\lambda_{m}$ and $\lambda_{e}$ are set to 0.6 and 80, respectively. While the Beta noise generally performs better than Gaussian noise in most settings, as shown in Table 2, there exist some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the synthetic noisy scores in Eq.(4), $\lambda_{m}$ for the combined denoising loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic noise generator to model-based one; 2) Over-amplifying these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components. Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise generator can further enhance the performance over the heuristic one.
@@ 5 Conclusion
6 Acknowledgement
This research was supported by the National Natural Science Foundation of China (U25A20445) and the advanced computing resources provided by the Supercomputing Center of the University of Science and Technology of China (USTC). This work was also supported by Kuaishou Technology, where author Wenyu Mao conducted this research during an internship.
@@ 6 Acknowledgement
All the results in this work are reproducible. We have discussed the optimal hyperparameters and the details on devices and software environments in Section 4.1 and Appendix D.2. For reproducibility, we release the code at https://github.com/maowenyu-11/DNR
@@ B.1 Derivation of the Three Objectives in our Solution
In this section, we illustrate the derivation of the Eq.(5). Similar to the variational Bayesian inference Kingma and Welling (2014), we start by analyzing the KL divergence between the real retriever scores and synthetic noisy scores:
@@
Note that a key difference between Eq.(8) and the conventional derivation in Kingma and Welling (2014) is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise generator that mimics the behavior of retriever scores. In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the real retriever scores $\mathbf{x}_{u}$ are not sampled by a manually defined distribution. Instead, an existing retriever $p_{\text{x}}$ serves as the prior distribution.
@@ D.3 Baselines
  • SASRec Kang and McAuley (2018) proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
  • Caser Tang and Wang (2018) introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
@@ E.2 Visualization of Learned Noises
Figure 5: The visualization of generated noise from different variants.
@@ E.5 Experiments on Larger Reranker Backbones
DNR is a backbone-agnostic framework for reranking tasks, utilizing and denoising the noisy scores from the previous retrieval stage through adversarial learning. Recently, LLMs have become a powerful tool for reranking tasks Mao et al. (2025b). However, text-based LLM4Rerank is not comparable in our setting offline or online, due to the lack of rich textual information (such as detailed item descriptions) that LLM-based rerankers fundamentally rely on. To evaluate the effectiveness of DNR in large models, we scaled up our PRM baseline by significantly increasing the number of its Transformer layers, serving as a “larger model” for comparison. We then tested our DNR framework, using this scaled-up PRM as its reranker backbone. The results are presented in the table 10.
@@ E.6 Online A/B Test
Detailed Experimental Settings for Online Test The multi-stage pipeline of the online system consists of three stages: retrieval, ranking, and reranking. In the final reranking stage, the top-6 items for user exposure are selected from a candidate set of 50 items generated by the ranking stage. During the ranking process, a scoring model computes multi-dimensional features for each item. Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$, which corresponds to the “realshow” metric in the online system. Note that the final reranking stage considers the combination of the retrieval stage and the middle ranking stage as a holistic previous stage, so the output PCTR scores from the ranking stage are considered as “retrieval score” in this final stage. DNR is applied to generate a new set of scores for 50 selected items and the top-6 items are chosen as the final exposure.
Experiment Results For our online experiments, the total traffic is randomly divided, with 12.5% allocated to the DNR method and 12.5% to the baseline method. Each experiment is conducted for at least 7 days to ensure the reliability of the results. The DNR method demonstrates a significant improvement in realshow(, i.e., the cumulative number of videos watched by users), increasing it by 1.089%. These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow. Additionally, the industrial recommendation environment involves various other metrics in addition to the realshow metric, including but not limited to app usage time, watch time, share-rate, like-rate, and comment-rate. We provide the corresponding results in Table 11. We can see that, in our scenario, watch-time and share-rate are slightly negatively impacted, while like-rate and comment-rate are slightly positively impacted, indicating a potential trade-off between metrics. Notably, while the watch time exhibited a slight decrease, we additionally observed that the long-term DAU slightly increased by 0.01% [-0.012%,+0.012%], which is nearly reaching a statistically significant improvement. This potentially suggests that the slight reduction in viewing duration is likely attributable to users exploring a broader range of clicked videos, while more significant long-term indicators remain positive. Beyond these, the impact on all other metrics is not statistically significant, which validates the effectiveness of our method. It is worth noting that achieving performance gains across all online metrics is inherently a multi-objective challenge—an area that remains a key focus for future work. Additionally, we believe applying DNR to other xTR signals (e.g., watch time and share-rate) holds significant potential for further improvements.
@@
E.7 Experiments on Diverse Retrievers
To better illustrate the generalizability of DNR across different retrievers, we implemented the dual-tower (DT) model as a stronger retriever. The comparison results in the retrieval stage are summarized in Table 12, and the comparison results in the reranking stage are summarized in Table 13. We can see that DT+DNR outperforms DT+PRM, validating the generalizability of our method and the sustained advantage of the denoising framework.
Theoretically, there will always be an empirical error that upper bounds the overall recommender system, which aims to minimize the "optimizable error", and there will be no room for the reranker to "denoise the scores" if the retriever itself achieve this bound. Yet, this theoretical bound is still impractical for a single retriever in most industrial platforms, due to various reasons including (but not limited to) the computational bottleneck, multi-task trade-off, and filtering on large-scale dynamic candidate pool. A more advanced retriever means a smaller remaining error that can be optimized, which indicates that the reranker will solve a simpler denoising problem and simpler reranker design. In this case, a diminishing marginal gain only indicates a diminishing computational requirement in the reranking stage. In general, the advantage of the denoising formulation is maintained regardless of the retriever employed.
Table 12: The AUC performance of the different retrievers, including the matrix factorization (MF) model and dual-tower model.
ML-1M Kuaivideo Book
MF 79.20% 89.62% 88.82%
DT 79.86% 90.38% 90.0%
Table 13: The performance comparison of rerankers on different retrievers.
Methods ML-1M Kuaivideo Book
NDCG@6 MAP@6 NDCG@6 MAP@6 NDCG@6 MAP@6
MF+PRM 72.85 62.21 55.93 42.97 76.88 66.44
MF+PIER 75.99 65.98 65.11 52.55 80.22 71.62
MF+DNR 77.67 68.00 70.15 58.12 83.53 75.54
DT+PRM 76.68 67.37 65.77 53.41